OpenStudy (anonymous):

The slope of a line is 12 and the point (1, 10) is on the line. Express the line in slope-intercept form

OpenStudy (amistre64):

y - y1 = m(x-x1) ; where P(x1,y1) = (1,10), and m= slope y - 10 = 12(x-1) y -10 = 12x -12 y = 12x -12+10 y = 12x -2

OpenStudy (anonymous):

The slope of a line is 13 and the point (3, 19) is on the line. Express the line in slope-intercept form.

OpenStudy (anonymous):

same question, just values are different:)